OpenStudy (anonymous):

Two numbers, x and y, are randomly chosen from the interval [0,1]. What is the expected value of (x+y)?

OpenStudy (anonymous):

what is the expected value of \(x\) ?

OpenStudy (anonymous):

it is clear that the expected value of \(x\) is \(\frac{1}{2}\) ?

OpenStudy (anonymous):

not really?

OpenStudy (anonymous):

well, one thing we could say, what else would it be?

OpenStudy (anonymous):

you pick a number at random in the interval, why should any number be favored? have of the length of the interval is one half do you have a definition for the expected value of a continuous variable?

OpenStudy (anonymous):

not really no

OpenStudy (anonymous):

hmm you have to have something right?

OpenStudy (anonymous):

oh wait. I see. since they are all equally chosen, then there is 1(the distance) divided by 2 (the # of variables). Yep! makes sense

OpenStudy (anonymous):

...right?

OpenStudy (anonymous):

wait. then y and x would both be 1/2 so 1/2+1/2 = 1 Ooh! the answer is 1! Thanks!

OpenStudy (anonymous):

hold on for a second, i don't want to steer you wrong

OpenStudy (anonymous):

darn

OpenStudy (anonymous):

ok i just had to check you are right it is \(\frac{1}{2}+\frac{1}{2}=1\) for your answer

OpenStudy (anonymous):

yay! is my reasoning right or...?

OpenStudy (anonymous):

there is an actual definition in terms of integrals, but i have a feeling you have not covered it yes, reasoning is right

OpenStudy (anonymous):

Thanks so much :D

OpenStudy (anonymous):

well actually it has nothing to do with the number of variables the expected value of \(x\) is \(\frac{1}{2}\) as is the expected value of \(y\) so since they are both one half, the expected value of the sum is one

OpenStudy (anonymous):

oh. ok.

OpenStudy (anonymous):

the reason the expected value is \(\frac{1}{2}\) is because \[\int_0^1xdx=\frac{1}{2}\] but i sort of doubt you have this in your class i could be wrong

OpenStudy (anonymous):

yah. youre right. I have no idea what you mean. still a newb XD

OpenStudy (anonymous):

don't worry about it